OpenStudy (anonymous):

Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (amistre64):

might want to break it down into right triangles

OpenStudy (amistre64):

or, create component vectors to work with ....

OpenStudy (anonymous):

direction of force is in the second quadrant using the law of parallelogram for vector addition i.e resultant force=

OpenStudy (anonymous):

\[\sqrt{f1^{2}+f2^{2}-2f1f2\cos{\theta}}\]

OpenStudy (anonymous):

where theta is the angly between the two forces

OpenStudy (anonymous):

238 pounds

OpenStudy (dumbcow):

i would split it into horizontal and vertical components \[R_x = 200\cos 60 + 100\cos 170\] \[R_y = 200\sin 60 + 100 \sin 170\] magnitude: \[|R| = \sqrt{R_x^{2} + R_y ^{2}}\] direction: \[\theta = \tan^{-1} (\frac{R_y}{R_x})\]